A master painter can paint a house in 12 hours. An assistant can paint the house in 16 hours. How long will it take the 2 painters to do the job if they work together to paint the house?
Found 2 solutions by addingup, ikleyn:
Answer by addingup(3677)

1/12+1/16 = 1/t
multiply all times 192t
16t+12t = 192
28t = 192
t = 6.86 hour if they work together

Answer by ikleyn(52806)

.
You can find a variety of solved joint-work problems in the lessons

    - Using Fractions to solve word problems on joint work 

    - Solving more complicated word problems on joint work 

    - Selected joint-work word problems from the archive 

in this site.